This website uses cookies. To read more about our use of cookies, click here.
To not be bothered by this message again, click the OK button.

ART Staff Blog Feed


November Progress Update

posted by Ed on 30 November 2016

After last months brief period of observations we are sorry to report that the ART is still unable to perform autonomous observations due to continuing hardware faults.
Read more »


PIRATE data contributes to Gaia follow up

posted by Ed on 1 November 2016

This week's Gaia image of the week is a graph of a microlensing event, which many observatories around the world have been helping to follow up. Included among these is our newly commissioned PIRATE telescope, seen credited at the bottom.


October Progress Update

posted by Ed on 13 October 2016

Recently, after a successful service trip to Tenerife, some users with outstanding jobs in the ART queue had images returned. Over the nights of the 1st, 2nd and 3rd of October approximately 300 images were taken autonomously.
Read more »


September Progress Update

posted by Ed on 12 September 2016

Work continues on trying to re-establish observations with the Autonomous Robotic Telescope (ART). On the night of Thursday 8th September a test run was carried out and a number of jobs were completed with Galaxy Cam. The images returned appeared to be of good quality, however further work is required on the alignment system of the telescope.

We hope to establish a fuller service soon and a service visit is scheduled for the end of September to aid in this effort.
Read more »


August Progress Update

posted by Ed on 28 July 2016

A team have visited the Teide Observatory to perform maintenance on the Autonomous Robotic Telescope (ART) and oversee the construction of the new facilities that will form the OpenScience Observatories.
Read more »


Welcome back!

posted by Ed on 21 June 2016

Welcome back to The website and the telescope have now completed the transfer to the Open University.

If you are a registered user who agreed to have your account transferred, you should find that your archive is as you left it.

If you have any problems with the use of the website or you account then please contact the team.
Read more »


Lunar Eclipse

posted by Chris on 1 October 2015


On the night of Sunday 27th to Monday 28th September 2015 a total lunar eclipse coincided with a "Supermoon" (a full Moon at the same time as the Moon being at its closest point to the Earth). With some adaptations to our scheduling software, we captured the following images.

Read more »


Tenerife summer maintenance 2015 update 2

posted by Dan on 13 August 2015

On reviewing last night's alignment run data we found that the quality of the mapped points was insufficient, and would not be usable to create an acceptable pointing model.

Today we have returned to the observatory with the aim of carrying out another alignment run this evening. In preparation for this second alignment run we have adjusted a number of elements of the mount RA drive, in the hope that we will improve the pointing and tracking performance in this axis. We are cautiously optimistic that we have made a breakthrough in this area.. but we won't know for sure until we review the full data set generated by this evening's alignment run.


Tenerife summer maintenance 2015 update

posted by Dan on 29 July 2015

The summer maintenance trip is now well underway. We have completed the work to re-grease the mount, and to run replacement wiring for the Galaxy camera focuser.

We are currently aiming to bring the system back online shortly with the Galaxy focuser working without its temperature sensor. Once we have the system back online we will be able to continue the work to restore the focuser functionality during the day, minimising the interruption to the operation of the system. To this end we are planning to perform a mapping run tonight with the hope of bringing the system back up tomorrow evening.


Tenerife summer maintenance 2015: Focuser repair

posted by Dan on 26 July 2015

The delayed 2015 summer maintenance trip has started today. Our focus on this trip will be to repair the faulty focuser on the galaxy camera.

Our initial investigations today have revealed that a fault with the wiring running through the mount has caused the focuser unit to fail. We have run tests with cables bypassing the mount wiring and found that the focus position changes correctly, however even with replacement wiring the temperature sensor is not responding.

We have separately tested the temperature sensor and verified its operation, leading us to believe that there is a fault with the temperature sensor input on the focuser control unit.

We are currently working on a number of solutions to this problem.. but we'll tell you about them later. For now I can say that we're in the process of replacing the faulty wiring and are hoping to re-grease the mount (a part of the regular summer maintenance work) tomorrow. If all goes to plan we will be realigning the system in a few days time, allowing us to return to operation with the cluster and constellation cameras as we work to complete the repairs on the galaxy focuser unit.

© 2017 The Open University